Beryl, Ena and Dave are on a coach trip and they stop at a cafe to get a cake or two. Beryl eats 3 cakes and Dave eats 2 more cakes than Ena.
Write an expression for the total number of cakes the three friends ate.

Answer by RAY100(1637)

let E= no of cakes Ena ate
let 3 = no of cakes Beryl ate
let E+2= no of cakes Dave ate
Total=E+3+E+2=2E+5